OpenStudy (anonymous):

Find the equation of the tangent line to the equation y=x^2 at x=1. Then, this line is also tangent to a circle with center (8, 0) at x = 2. Find the equation of this circle. I got y=2x+1 for my tangent line (which is wrong), and had no idea how to get the second part.

OpenStudy (anonymous):

So when they say find tangent line to y=x^2 that means take derivative of that first.

OpenStudy (anonymous):

Do you know derivative of x^2 ?

OpenStudy (anonymous):

yes, it is 2x. and then plug in x=1 and you get 2. therefore the slope is 2 and then plug x=1 back in to the original equation and get 1. therefore y=2x+1.....that was how I got it at least

OpenStudy (anonymous):

yet when I submitted that for my answer, it said it was incorrect...

OpenStudy (campbell_st):

so next find the perpedicular distance from a point to a line the line is y = 2x + 1 the point is (8, 0) this will give the radius of the circle. Also the tangent to a circle is perpendicular to the radius. then the standard form is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre and r is the radius

OpenStudy (anonymous):

y=2x-1 should be the correct tangent equation.

OpenStudy (anonymous):

why is it -? the equation is y=mx+b and there isn't a negative value to make it negative...

OpenStudy (anonymous):

Graph x^2. at x=1 the tangent line is represented by 2x. So if you think of rise of run. Imagine going to the left from x=1 on the graph, and you should go down two units.

OpenStudy (campbell_st):

the y intercept is -1 the point on the parabola is (1, 1) the general from is y = mx + b so y = 2x + b substitute the point x = 1 and y = 1 to find b...

OpenStudy (anonymous):

So if the y intercept is -1 wouldn't the equation be y=2x-1?

OpenStudy (campbell_st):

that's correct...

OpenStudy (campbell_st):

so you need to now find the perpendicular distance from a line to a point to get the radius of the circle.

OpenStudy (anonymous):

ok, I get that now...now what about the second part. I know campbell_st started it, but is there anything else I should know before I try to do it?

OpenStudy (anonymous):

Just use the distance formula between the two points.

OpenStudy (campbell_st):

nope the tangent and radius of the circle are perpendicular to each other. so find the radius, you know the centre so substitute it into the standard form of the equation.

OpenStudy (anonymous):

ok, and just to check, the radius is 6 since the center is at (8,0) and and they want to know for the point when x=2 (8-2=6)

OpenStudy (campbell_st):

so if you have a line \[ax + by + c = 0\] the perpendicular distance is \[d = \frac{\left| ax_{1} + by_{1} + c \right|}{\sqrt{a^2 + b^2}}\] so you have \[d = \frac{\left| 2\times 8 - 1\times0 -1 \right|}{\sqrt{2^2 + 1^2}}\]

OpenStudy (anonymous):

assuming that is true, so far, for my equation, I have 36+y^x=36 which does not seem right

OpenStudy (anonymous):

I have no clue what you just typed and have no idea what I am supposed to do with that...

OpenStudy (campbell_st):

well a slight correction, at x = 2 is a point on the tangent line and the circle's circumference... a point of intersection.... so if the tangent is y = 2x - 1 when x = 2 the point is y = 2*2 - 1 so the point is (2, 3)

OpenStudy (anonymous):

I am honestly so lost. Can someone just do the work and show me because I have no clue what to do at all

OpenStudy (campbell_st):

ok... it's a formula for finding the perpendicular distance from a point to a line

OpenStudy (anonymous):

ok, but I honestly need you to tell me what to do because I am lost

OpenStudy (campbell_st):

ok... so you have 2 points (2, 3) and (8, 0) find the distance between them. this will be the radius and I've attached a graph

OpenStudy (campbell_st):

so leave the radius as a radical and then the standard form of a circle is \[(x - h)^2 + (y - k)^2 = r^2\] you know the centre h = 8, k = 0 and you calculate the radius... so substitute them

OpenStudy (anonymous):

what is the distance formula again?

OpenStudy (anonymous):

nevermind, I looked it up...the distance is root(45)

OpenStudy (campbell_st):

given 2 points \[(x_{1}, y_{1})~~~(x_{2},y_{2})\] the distance between them is \[d = \sqrt{x_{2} - x_{1})^2 + (y_{2} - y_{1})^2}\] or just use pythagoras' theorem 8 to 2 = 6 and up 3 so d^2 = 6^2 + 3^2

OpenStudy (anonymous):

i know this goes against what the purpose of this site is, and I do appreciate all of your help, but my homework is due soon, midnight my time, as it is online homework, can you please give me the answer as I still have other problems I need to do too. I have tried to do them while working on this one, but this has taken too much to focus on the others also. I am sorry, but I really need the answer. I will be willing to work it out later though

OpenStudy (campbell_st):

ok... so here is the guts of it  just find r^2. the circle is and substitute your value for r^2 \[(x - 8)^2 + y^2 = r^2\]

OpenStudy (anonymous):

r^2=45

OpenStudy (campbell_st):

great... now substitute it for the equation of the circle

OpenStudy (anonymous):

(x-8)+y^2=45 Thanks!

OpenStudy (campbell_st):

its (x - 8)^2

OpenStudy (anonymous):

I meant to type that...whoops!